OpenStudy (anonymous):

find two complementary angles in the ratio of 7:2

OpenStudy (anonymous):

7+2=9 90/9=10 10*2=20 10*7=70 20,70

OpenStudy (anonymous):

why do you add 7 and 2 ?

OpenStudy (anonymous):

Ok, let's do it algebraically, complemetry angles add up to 90 two angle are 7x, 2x 7x+2x=90 9x=90 x=10 7x=70 2x=20

hero (hero):

Some things just aren't very obvious

OpenStudy (anonymous):

lol or im just dumb.

hero (hero):

No....I've never seen this question before either, so I'm scratching my head too

hero (hero):

I think I get it though. As long as the ratio is 7/2, were good, so 70/20 works. And we know they're complementary, so they also have to add to 90. So 70 + 20 works.